OpenStudy (anonymous):

Question about vectors? * (v and w are vectors) Given v= (4, 2) and w= (-3, 5), find the coordinates of v + w. a. (1, 7) b. (1, -3) c. (7, 7) d. (7, -3) I got answer B. Is that correct?

OpenStudy (turingtest):

no it's not add each component seperately

OpenStudy (turingtest):

v=<a, b> u=<c, d> v+u=<a+c, b+d>

OpenStudy (anonymous):

That still gives me the same answer, though

OpenStudy (turingtest):

show your work please

OpenStudy (anonymous):

v= (4, 2) 4 is a, 2 is b w= (-3, -5) -3 is c, -5 is d v + w = (4 + -3), (2 + -5) v+w = (1, -3)

OpenStudy (turingtest):

the way you typed the problem above you have w=(-3,5), so which is it?

OpenStudy (anonymous):

I'm sorry, it's (-3, -5).

OpenStudy (turingtest):

then you were right all along :)

OpenStudy (anonymous):

Haha, thank you. Sorry about that!

OpenStudy (turingtest):

No worries, it happens to the best of us ;)